OpenStudy (anonymous):

I been stuck on this question forever so pls help me someone! is math We have to write the slope-intercept for the equation? this is the problem through: (3,2), parallel to y=x+2 i know that it starts like y-2= but then is it 1/1(fraction? im confused @satellite73 @misty1212 @inkyvoyd

OpenStudy (madhu.mukherjee.946):

when a line is parallel to another line then both of them have equal slopes

OpenStudy (anonymous):

okey so it is 1/1 right since theres not a fraction @madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

yeah

OpenStudy (phi):

parallel to y=x+2 you match that equation with the form y=mx+b y= x+2 y= mx+b it looks like b matches with 2, so 2 is the y-intercept what is the slope "m" ? if you see x by itself, you can think of it as 1*x (one times x) now it is more clear: y=1x+2 y=mx+b the slope m matches with 1 so the slope is 1 now you can use the point slope formula y-y0= m(x-x0) where m is the slope and (x0,y0) is a point in this problem, m is 1 and the point is (3,2) you write: y-2 =1(x-3) 1 times (x-3) is (x-3) (or you could "distribute" the 1 to get 1*x - 1*3 which is x-3 either way you get y-2=x-3 now add +2 to both sides to get y=x-3+2 or y= x-1 as the equation in slope-intercept form